# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v1

@@ 2.2 Verifiable Homomorphic Encryption
Verifiable Homomorphic Encryption is stated as follows.
@@
A Verifiable Homomorphic Encryption is a tuple of PPT algorithms
$\Pi=(\mathsf{KeyGen,Enc_{key},Eval_{evk},Verify_{sk},Dec_{sk}})$
@@ A Verifiable Homomorphic Encryption is a tuple of PPT algorithms
  • $(\textsf{pk}{},\textsf{sk}{})\leftarrow{\sf KeyGen}(1^{\lambda},f)$:
  • $(\textsf{ct}{}_{x},\tau_{x})\leftarrow{\sf Enc_{key}}(x)$ for ${\sf key}=\textsf{sk}{}$ or pk.
  • $(\textsf{ct}{}_{y},\tau_{y})\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$
  • $b\leftarrow{\sf Verify_{\textsf{sk}}{}}(\textsf{ct}{}_{y},\tau_{x},\tau_{y})$, the client accepts if $b=0$ and rejects if $b=1$.
@@
Definition 4 (Completeness).
A Verifiable Homomorphic Encryption scheme $\Pi$ is complete if the client always accepts a correct output.
$\Pr\left[0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y},\tau_{x},\tau_{y}):\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},f))\\ (\textsf{ct}{}_{x},\tau_{x})&\leftarrow\textsf{Enc}_{\mathsf{key}}(x)\\ (\textsf{ct}{}_{y},\tau_{y})&\leftarrow\textsf{Eval}_{\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]=1$
2.2.1 Adversary Model.
Before defining the security of a VHE scheme, we need to formalize the adversary model.
Types of Adversary.
In this paper, we assume two types of adversary: Malicious one and Covert one. The covert adversary is an adversary that executes attacks when the probability of detection is negligible, as assumed in [CKP+24]. For more details on the covert adversary, see [AL07].
Access to Oracles.
[…]
2.2.2 Security and Attack.
Now we define the security of VHE. This definition is a modification of the soundness notion presented in [VKH23]. Here, we assume that $\Pi$ is correct and complete.
Definition 5 (Security against Malicious Adversary).
A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of malicious adversaries if the client rejects an incorrect output with overwhelming probability in $\lambda$ for any malicious PPT adversary $\mathcal{A}_{m}$.
$\Pr\left[\begin{aligned} 0\leftarrow&{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y},\tau_{x},\tau_{y})\\ &\textsf{Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y})\neq f(x)\end{aligned}:\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},f))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{key}}(x)\\ (\textsf{ct}{}_{y},\tau_{y})&\leftarrow\mathcal{A}^{\textsf{evk}}_{m}(\textsf{ct}{}_{x},\tau_{x})\\ \end{aligned}\right]\leq 2^{-\lambda}$
Definition 6 (Security against Covert Adversary).
A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of covert adversaries if the client rejects an incorrect output with overwhelming probability in any covert PPT adversary $\mathcal{A}_{c}$.
[…]
@@ 3 Cryptanalysis on vADDG
In [ADDG24], Albrecht et al. presented an Oblivious PRF candidate and its extension to a verifiable OPRF, known as the ADDG and vADDG schemes, respectively. The ADDG scheme utilizes TFHE to construct an OPRF candidate: the client outsources homomorphic encryption of $\mathbf{x}$, and the server homomorphically computes $F(\mathbf{k},\mathbf{x})$ and returns it to the client. In this process, the OPRF protocol requires that the server does not learn either $\mathbf{x}$ or $\mathbf{z}$.
@@
First, the client inputs an encryption of a 128-bit string $\mathbf{x}\in\mathbb{Z}_{2}^{128}$. Then, using only homomorphic additions and constant multiplications without bootstrapping, the server homomorphically evaluates $F(\mathbf{k},\mathbf{x})$ on a 256-bit string $\mathbf{x}\in\mathbb{Z}_{2}^{256}$ using encryption of the server’s PRF key $\mathbf{k}$ Next, the server performs ${\sf CPPBS}_{(2,3)}$, a programmable bootstrapping with functionalities
@@
to obtain an encryption of $\mathbf{y}\in\mathbb{Z}_{3}^{256}$. Finally, again using only homomorphic additions and constant multiplications, the server computes an encryption of an 82-trit string $\mathbf{z}\in\mathbb{Z}_{3}^{82}$, which is the output of the homomorphic evaluation of the PRF. A notable feature of the ADDG scheme is that it requires only a bootstrapping depth of one: The key-switching key that converts the changed secret key after bootstrapping back to the original secret key is not provided. Thus, it prevents any further bootstrapping execution.
@@
    First, the server publishes verification values
    =$\mathfrak{R}=\{(\mathbf{x}^{\star}_{k},\mathbf{z}^{\star}_{k}=F(\mathbf{k},\mathbf{x}^{\star}_{k}))\}_{1\leq k\leq\kappa}$
@@
    where $\mathbf{x}^{\star}_{k_{j}}\overset{\$}{\leftarrow}\mathfrak{R}$ and $\gamma=\alpha\nu+\beta$.
@@
    Client verifies whether the following is correct.
@@ 3.2 Attack on vADDG
The security of the vADDG scheme relies on the hardness of constructing a deep cheating circuit under restricted depth [CHLR18], and it prevents forgery attacks by limiting the bootstrapping to only one. However, despite the bootstrapping depth limitation, there exists a cheating circuit that can forge the result with a successful attack probability for $(\alpha,\beta,\nu)=(105,10,11)$. The idea behind this attack is to create a pseudo-random characteristic function and implement it homomorphically. To build such a function, we utilize the function $\mathsf{Mult}^{k}_{(2,3)}:\mathbb{Z}_{2}^{k}\to\mathbb{Z}_{3}$ defined by
$\mathsf{Mult}^{k}_{(2,3)}(x_{1},\dots,x_{k})=\begin{cases}1&\text{if }x_{i}=1\text{ for all }i,\\[2.84526pt] 0&\text{otherwise.}\end{cases}$
First, the adversarial server receives the inputs $\textsf{ct}{}_{s}:=[\mathbf{x}_{s}]_{2}$ for all $1\leq s\leq\gamma$ where each $\mathbf{x}_{s}\in\mathbb{Z}_{2}^{128}$ is a 128-bit string. Then it evaluates $\mathsf{Mult}^{k}_{(2,3)}$ on $\textsf{ct}{}_{1},\textsf{ct}{}_{s},\dots,\textsf{ct}{}_{\gamma}$. If the TFHE parameters support programmable bootstrapping on $k$ input ciphertexts, TFHE programmable bootstrapping can evaluate ${\sf Mult^{k}_{(2,3)}}$ with only depth-one bootstrapping using encrypted table lookup.
With this programmable bootstrapping, we design a homomorphic characteristic function $\chi:\mathcal{C}^{128}\to\mathcal{C},\textsf{ct}{}_{s}\mapsto\textsf{ct}{}^{\mathsf{Mult}}_{s,u}$ as follows: First, choose distinct $k$ indices $u=\{u_{1},\dots,u_{k}\}\subset\{1,\dots,128\}$. Next, for each string $\textsf{ct}{}_{s}$, evaluate
$\textsf{ct}{}^{\mathsf{Mult}}_{s,u}:=\textsf{Eval}_{{\sf Mult}^{k}_{(2,3)}}(\textsf{ct}{}_{s,u_{1}},\dots,\textsf{ct}{}_{s,u_{k}}).$
Then, this characteristic function $\chi:\textsf{ct}{}_{s}\mapsto\textsf{ct}{}^{\mathsf{Mult}}_{s,u}$ is a homomorphic pseudo-random function which outputs $[1]_{3}$ with approximate probability $(1/2)^{k}$, otherwise $[0]_{3}$. With these $\gamma$ outputs $(\textsf{ct}{}^{\mathsf{Mult}}_{s,u})_{1\leq s\leq\gamma}$, the adversary can try to forge the PRF $F$ by adding $(\textsf{ct}{}^{\mathsf{Mult}}_{s,u})_{1\leq s\leq\gamma}$ to the output strings $\left(\textsf{Eval}_{F(\mathbf{k,-})}(\textsf{ct}{}_{s})\right)_{1\leq s\leq\gamma}$. Note that $\chi(\textsf{ct}{}_{s})$ is the same for the same inputs and so this forgery is not detected if $\textsf{ct}{}^{\mathsf{Mult}}_{s,u}$ is 0 for the $\beta$ verification values. The success probability is given in Theorem 1.
@@
1:procedure $\mathsf{vADDG\_Attack}$($(\textsf{ct}{}_{s}),k$) $\triangleright$ $\textsf{ct}{}_{s}=[\mathbf{x}_{s}]_{2}=([x_{s,1}]_{2},\dots,[x_{s,128}]_{2})$ for $1\leq s\leq\gamma$.
2:     $u_{1},\dots,u_{k}\leftarrow\{1,\dots,128\}$
3:     $\mathbf{t}\leftarrow\mathbb{Z}_{3}^{82}$ for $\mathbf{t}\neq({0,0,\dots,0})$triangleright Select the trits to forge.
4:     for $s=1$ to $\gamma$ do
5:         $\textsf{ct}{}_{s,u}^{\sf Mult}\leftarrow\textsf{Eval}_{\mathsf{Mult}^{k}_{(2,3)}}([x_{s,u_{1}}]_{2},\dots,[x_{s,u_{k}}]_{2})$
6:         $[\mathbf{z}_{s}]_{3}\leftarrow\textsf{Eval}_{F(\bf k,-)}(\textsf{ct}{}_{s})$triangleright $[\mathbf{z}_{s}]_{3}=([z_{s,1}]_{3},\dots,[z_{s,82}]_{3})$.
7:         $\textsf{ct}{}^{\sf forged}_{s}\leftarrow[{\bf z}_{s}]_{3}+\textsf{ct}{}_{s,u}^{\sf Mult}\cdot\mathbf{t}$
8:     end for
9:     return $\left(\textsf{ct}{}_{s}^{\sf forged}\right)_{1\leq s\leq\gamma}$
10:end procedure
@@
Assume that the TFHE scheme supports $k$ inputs per one programmable bootstrapping. Then the expected probability of the depth one circuit in Algorithm 1 successfully forging the output is
$\left(1-\left(1-2^{-k}\right)^{\alpha}\right)\cdot\left(1-2^{-k}\right)^{\beta}.$
@@
The forgery succeeds when $\textsf{ct}{}_{s,u}^{\sf Mult}=[0]_{3}$ for all $s$ correspond to the $\beta$ verification values and at least one $\textsf{ct}{}_{s,u}^{\sf Mult}=[1]_{3}$ for $s$ corresponds to $\alpha$ message values. Let $p$ be the probability that $\textsf{ct}{}_{s,u}^{\sf Mult}$ became an encryption of 1. In the Algorithm 1, the expected value of $p$ over all input values is $2^{-k}$. The attack success probability is obtained by subtracting the probability that all values become zero, $\left(1-p\right)^{\alpha+\beta}$, from the probability that all $s$ corresponding to $\beta$ become zero, $(1-p)^{\beta}$. Thus, the success probability is given by $Q(p):=\left(1-p\right)^{\beta}-\left(1-p\right)^{\alpha+\beta}.$
Substituting $p=2^{-k}$, we have the desired result
$Q(2^{-k})=\left(1-\left(1-2^{-k}\right)^{\alpha}\right)\cdot\left(1-2^{-k}\right)^{\beta}.$
For example $k=5$, with the parameter $(\alpha,\beta,\nu)=(105,10,11)$, $Q(2^{-k})\cong 70.2\%.$
@@
The vADDG scheme in the parameter $(\alpha,\beta,\nu)=(105,10,11)$ is insecure against the presence of $\mathcal{A}_{m}$.
@@
The success probability of this attack does not change with an increase in the parameter $\nu$. Also, the probability does not change significantly with a decrease in the parameter $\alpha$, since $Q(p)=(1-(1-p)^{\alpha})(1-p)^{\beta}$ has a lower bound of $p(1-p)^{\beta}$ and an upper bound of $(1-p)^{\beta}$, which are both independent of $\alpha$. Thus in our discussion of parameters, we assume that $\alpha$ is sufficiently large so that $Q(p)\lesssim(1-p)^{\beta}$. To bound the attack success probability $(1-p)^{\beta}$, two conditions are needed: a nonzero lower bound for $p$, and the value of $\beta$ large enough relative to this lower bound. The nonzero lower bound for $p$ that can be obtained with $k$ ciphertext inputs is $1/2^{k}$ via $\mathsf{Mult}^{k}_{(2,3)}$. Although the maximum of $k$ depends on specific TFHE parameters, if we roughly may assume $1\leq k\leq 8$ Also, $\beta$ must be large enough to match this bound.
In particular, we have $(1-2^{-k})^{\beta}\leq 2^{-\lambda}$ if the following condition holds: $\beta\geq\frac{\lambda\cdot 2^{k}}{\log e}$. Therefore, to target an 80-bit security with $k=8$, $\beta$ must be at least $\beta\geq 14196\approx\frac{80\cdot 2^{8}}{\log e}$. In particular, if $k$ is fixed and $\beta$ is sufficiently large, increasing $\alpha$ arbitrarily does not affect this attack.
In [ADDG24], other cheating strategies are mentioned with success probabilities of
$\frac{1}{\binom{\gamma}{\beta}}\quad\text{and}\quad\frac{\alpha}{\binom{\gamma}{\nu}}.$
By setting $\beta=14196$, we newly propose the following parameter set that satisfies 8080{80}80-bit security:
$(\alpha,\beta,\nu)=(136687,14196,5).$
This parameter has an overhead size of approximately ×$\times 5.10$ but requires a large set of inputs.

### v2

@@ 2.2 Verifiable Homomorphic Encryption
Verifiable Homomorphic Encryption is stated as follows. Many works in the literature introduce a tag as part of the output of Enc and Eval, but we will treat it as part of the ciphertext, when it exists. Let $\mathcal{M}$ be a plaintext space and let $\mathcal{F}\subset\{f_{\alpha}\mid f_{\alpha}:\mathcal{M}^{n_{\alpha}}\to\mathcal{M}\}$ be a set of functions on tuples of $\mathcal{M}$.
@@
A Verifiable Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
$\Pi=(\mathsf{KeyGen,Enc,Eval,Verify,Dec})$
@@ A Verifiable Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
  • $\textsf{pk}{},\textsf{sk}{}\leftarrow{\sf KeyGen}(1^{\lambda},\mathcal{F})$:
  • $\textsf{ct}{}_{x}\leftarrow{\sf Enc_{pk}}(x)$ for $x\in\mathcal{M}$.
  • $\textsf{ct}{}_{y}\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$ where $f\in\mathcal{F}$.
  • $b\leftarrow{\sf Verify_{\textsf{sk}}{}}(\textsf{ct}{}_{y})$, the client accepts if $b=0$ and rejects if $b=1$.
@@
Definition 4 ($\delta$-Completeness).
A Verifiable Homomorphic Encryption scheme $\Pi$ is $\delta$-complete if the client always accepts a correct output.
$\Pr\left[0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y}):\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta$
2.2.1 Malicious Adversary Model.
We now formally define the security of VHE. We assume that the underlying scheme $\Pi$ satisfies both $\delta$-correctness and $\delta$-completeness for some negligible $\delta$. Throughout this work, we consider two types of adversaries: malicious and covert.
Definition 5 (Security against Malicious Adversary, modified from [VKH23]).
A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of malicious adversaries if the client rejects an incorrect output with overwhelming probability in $\lambda$ for any malicious PPT adversary $\mathcal{A}_{\sf mal}$.
$\Pr\left[\begin{aligned} 0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y})\\ \textsf{Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y})\neq f(x)\end{aligned}:\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\mathcal{A}^{\textsf{evk}}_{\sf mal}(\textsf{ct}{}_{x})\\ \end{aligned}\right]\text{ is negligible.}$.
2.2.2 Covert Adversary.
Originally introduced in the context of secure multiparty computation in [AL07], the covert adversary model can be extended to the VHE setting, given that HE can be regarded as a special case of two-party computation, and considered in [CKP+24]. The $\varepsilon$-covert adversary model, or simply the covert adversary model, assumes that any adversarial deviation is detected with probability at least $\varepsilon$. A customary choice of the deterrence factor $\varepsilon$ is a noticeable probability ($1/{\sf poly}(\lambda)$).Unless otherwise stated, we use the term covert adversary to refer to a $1/{\sf poly}(\lambda)$-covert adversary. In this paper, we provide an informal definition of the covert adversary model. For a formal treatment, we refer the reader to the Definition 3.4 from [AL07].
Definition 6 (Security against Covert Adversaries, informal).
Let $\varepsilon\in[0,1]$. A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of $\varepsilon$-covert adversaries if for any PPT adversary $\mathcal{A}$, any deviation from the prescribed protocol is detected with probability at least $\varepsilon$.
2.2.3 Access to Oracles.
We assume that the adversary can access to the evaluation keys, but not to the decryption, verification, or encryption oracles. In the attack on REP and PE, we always assume that the adversary has rlk, and in the attack on vADDG, we assume that the adversary has $\sf btk$. The only exception is that in the case of PE, we assume that the adversary can access the ReQ protocol. We denote adversary models as follows: for example, $\mathcal{A}_{\sf mal}^{\textsf{rtk}{}_{1},\mathcal{O}_{\textsf{ReQ}}}$ denotes a malicious adversary who can access $\textsf{rtk}{}_{1}$, rlk, and ReQ, while $\mathcal{A}_{\sf cov}$ denotes an $1/{\sf poly}(\lambda)$-covert adversary who can access only rlk. (Since we always assume that all adversaries can access rlk, we omit it from the notation.)
@@ 3 Cryptanalysis on vADDG
In [ADDG24], Albrecht et al. presented an Oblivious PRF candidate and its extension to a verifiable OPRF, known as the ADDG and vADDG schemes, respectively. The ADDG scheme utilizes TFHE to construct an OPRF candidate: the client outsources homomorphic encryption of a bit string $\mathbf{x}$, and the server homomorphically computes $\bf z=F(\mathbf{k},\mathbf{x})$ with its PRF key $\bf k$ and returns it to the client. In this process, the OPRF protocol requires that the server does not learn either $\mathbf{x}$ or $\mathbf{z}$.
@@
First, the client inputs an encryption of a 128-bit string $\mathbf{x}\in\mathbb{Z}_{2}^{128}$. Then, using only homomorphic additions and constant multiplications without bootstrapping, the server homomorphically evaluates $\mathbf{y}\in\mathbb{Z}_{2}^{256}$ using the PRF key $\mathbf{k}$. Next, the server performs ${\sf CPPBS}_{(2,3)}$ (Circuit Privacy Programmable Bootstrapping), a programmable bootstrapping with functionalities
@@
to obtain an encryption of $\mathbf{y}$ over $\mathbb{Z}_{3}^{256}$. Finally, again using only homomorphic additions and constant multiplications, the server computes an encryption of an 82-trit string $\mathbf{z}\in\mathbb{Z}_{3}^{82}$, which is the output of the homomorphic evaluation of the PRF. A notable feature of the (v)ADDG scheme is that it consumes only a bootstrapping depth of one: By the removal of the key-switching key, this scheme prevents further bootstrapping executions after one bootstrapping.
$\boxed{\textsf{Eval}_{F_{\bf k}}:{[\mathbf{x}]_{2}^{\textsf{sk}}{}}\in\mathcal{C}^{128}{\mapsto}{[\mathbf{y}]_{2}^{\textsf{sk}}{}}\in\mathcal{C}^{256}\overset{{\sf CPPBS}_{(2,3)}}{\mapsto}{[\mathbf{y}]_{3}^{\textsf{sk}{}^{\prime}}}\in\mathcal{C}^{256}\mapsto{[\mathbf{z}]_{3}^{\textsf{sk}{}^{\prime}}}\in\mathcal{C}^{82}}$
Figure 1: Workflow of the Homomorphic PRF suggested in [ADDG24]. The superscripted $\mathsf{sk}$ denotes that the secret key is $\mathsf{sk}$.
@@
    First, the server publishes $\kappa$ verification values
    =$\mathfrak{R}=\{(\mathbf{x}^{\star}_{k},\mathbf{z}^{\star}_{k})\}_{1\leq k\leq\kappa}\quad(\mathbf{z}^{\star}_{k}=F(\mathbf{k},\mathbf{x}^{\star}_{k}))$
@@
    where $(\mathbf{x}^{\star}_{k_{j}},\mathbf{z}^{\star}_{k_{j}})\overset{\$}{\leftarrow}\mathfrak{R}$ and $\gamma=\alpha\nu+\beta$.
@@
    Client verifies whether the followings are correct.
@@ 3.2 Attack on vADDG
The vADDG scheme is designed to prevent forgery attacks by leveraging the assumed hardness of building deep circuits under depth restrictions [CHLR18], and by restricting bootstrapping to a single level. In [ADDG24], the authors propose a concrete attack strategy under this model and determine secure parameters to resist such attacks, leading to the choice of $(\alpha,\beta,\nu)=(105,10,11)$. However, we show that a successful forgery remains possible under this parameter set with non-negligible probability.
The idea behind this attack is to create a pseudo-random characteristic function and implement it homomorphically. To build such a function, we follow the following procedure. First, the adversarial server receives the inputs $\textsf{ct}{}_{s}:=[\mathbf{x}_{s}]_{2}$ for all $1\leq s\leq\gamma$ where each $\mathbf{x}_{s}\in\mathbb{Z}_{2}^{128}$ is a 128-bit string. We design a homomorphic characteristic function $\chi:\mathcal{C}^{128}\to\mathcal{C},\textsf{ct}{}_{s}\mapsto\overline{\textsf{ct}{}_{s,u}}$ as follows: First, Let an index $u\in\{1,\dots,128\}$ be given. Next, for each ciphertext string $\textsf{ct}{}_{s}$, evaluate
$\overline{\textsf{ct}{}_{s,u}}:={\sf CPPBS}_{(2,3)}(\textsf{ct}{}_{s,u}).$
Then, this characteristic function $\chi:\textsf{ct}{}_{s}\mapsto\overline{\textsf{ct}{}_{s,u}}$ is a homomorphic pseudo-random function which outputs $[1]_{3}$ with approximate probability $1/2$, otherwise $[0]_{3}$. With these $\gamma$ outputs $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$, the adversary can try to forge the PRF $F$ by adding $(\overline{\textsf{ct}{}_{s,u}})_{1\leq s\leq\gamma}$ to the output strings $\left(\textsf{Eval}_{F(\mathbf{k,-})}(\textsf{ct}{}_{s})\right)_{1\leq s\leq\gamma}$. For example, for ${\bf t}=(1,0,\dots,0)$, we can forge the PRF by evaluating
$\textsf{Eval}_{F({\bf k},-)}(\textsf{ct}{}_{s})+{\bf t}\cdot\chi(\textsf{ct}{}_{s}).$
Note that $\chi(\textsf{ct}{}_{s})$ is the same for the same inputs and so this forgery is not detected if $\chi({\textsf{ct}{}_{s}})$ is 0 for the all $\beta$ verification values. The success probability is given in Theorem 1.
@@
1:procedure $\mathsf{vADDG\_Attack}$($(\textsf{ct}{}_{s})_{1\leq s\leq\gamma},u$) $\triangleright$ $u\in\{1,\dots,128\}$ $\textsf{ct}{}_{s}=[\mathbf{x}_{s}]_{2}=([x_{s,1}]_{2},\dots,[x_{s,128}]_{2})$ for $1\leq s\leq\gamma$.
2:     $\mathbf{t}\leftarrow\mathbb{Z}_{3}^{82}$ for $\mathbf{t}\neq({0,0,\dots,0})$triangleright Select the trits to forge.
3:     for $s=1$ to $\gamma$ do
4:         $\overline{\textsf{ct}{}_{s,u}}\leftarrow{\sf CPPBS}_{(2,3)}([x_{s,u_{1}}]_{2})$
5:         $[\mathbf{z}_{s}]_{3}\leftarrow\textsf{Eval}_{F(\bf k,-)}(\textsf{ct}{}_{s})$triangleright $[\mathbf{z}_{s}]_{3}=([z_{s,1}]_{3},\dots,[z_{s,82}]_{3})$.
6:         $\textsf{ct}{}^{\sf forged}_{s}\leftarrow[{\bf z}_{s}]_{3}+\overline{\textsf{ct}{}_{s,u}}\cdot\mathbf{t}$
7:     end for
8:     return $\left(\textsf{ct}{}_{s}^{\sf forged}\right)_{1\leq s\leq\gamma}$
9:end procedure
@@
For a given parameter $(\alpha,\beta,\nu)$, the expected probability that the depth one circuit in Algorithm 1 successfully forges the output is
$\left(1-\left(1/2\right)^{\alpha}\right)\cdot\left(1/2\right)^{\beta}.$
@@
The forgery succeeds when $\overline{\textsf{ct}{}_{s,u}}=[0]_{3}$ for all $s$ correspond to the $\beta$ verification values and at least one $\overline{\textsf{ct}{}_{s,u}}=[1]_{3}$ for $s$ corresponds to $\alpha$ message values. Let $p$ be the probability that $\overline{\textsf{ct}{}_{s,u}}$ became an encryption of 1. In the Algorithm 1, the expected value of $p$ over all input values is $1/2$. The attack success probability is obtained by subtracting the probability that all values become zero, $\left(1-p\right)^{\alpha+\beta}$, from the probability that all $s$ corresponding to $\beta$ become zero, $(1-p)^{\beta}$. Thus, the success probability is given by $Q(p):=\left(1-p\right)^{\beta}-\left(1-p\right)^{\alpha+\beta}.$ Substituting $p=1/2$, we have the desired result $\left(1-\left(1/2\right)^{\alpha}\right)\cdot\left(1/2\right)^{\beta}.$
@@
The vADDG scheme with parameters $(\alpha,\beta,\nu)=(105,10,11)$ provides at most 10-bit security against a malicious adversary.
In the covert adversary setting, this forgery is detected if at least one of the $\beta$ verification values is altered, which occurs with probability $1-(1/2)^{\beta}$. Therefore, a covert adversary cannot successfully carry out this attack.
@@
The success probability of this attack does not change with an increase in the parameter $\nu$. Also, the probability does not change significantly with a decrease in the parameter $\alpha$, since $Q(p)=(1-(1/2)^{\alpha})(1/2)^{\beta}$ has a lower bound of $(1/2)\cdot(1/2)^{\beta}$ and an upper bound of $(1/2)^{\beta}$, which are both independent of $\alpha$. Thus in our discussion of parameters, we only concerns $\beta$. To ensure $80$-bit security, $\beta$ must exceed 80.
In [ADDG24], other cheating strategies are mentioned with success probabilities of $1/\binom{\gamma}{\beta}\text{ and }\alpha/\binom{\gamma}{\nu}.$ By setting $\beta=80$, we obtain the following two new parameter sets that satisfy 80-bit security:
$(\alpha,\beta,\nu)={(28440,80,6)}\text{, or }(189,80,10).$
The first parameter has an overhead size ($(\alpha\nu+\beta)/\alpha$) of approximately ×$\times 6.0$ but requires a larger set of inputs, meanwhile the second parameter has an overhead size of approximately ×$\times 10.4$ but requires a smaller set of inputs.